# A Simple, Optimal and Efficient Algorithm for Online Exp-Concave Optimization

### v1

@@ 2.2.2 Online Quasi-Newton Step (OQNS)
$\Phi_{t}(\mathbf{x})$ $\triangleq-\eta d\log(1-\left\|\mathbf{x}\right\|_{2}^{2})+\frac{d+\eta}{2}\left\|\mathbf{x}\right\|_{2}^{2}+\sum_{s=1}^{t}\left(\nabla f_{s}(\mathbf{x}_{s})^{\mathsf{T}}\mathbf{x}+\frac{\gamma}{2}\left(\nabla f_{s}(\mathbf{x}_{s})^{\mathsf{T}}(\mathbf{x}-\mathbf{x}_{s})\right)^{2}\right).$ (7b)
@@
Evaluating the Hessian-inverse still takes $O(d^{\omega})$ time and OQNS mitigates this issue by approximating the Hessian-inverse gradient product with incremental updates. The approximation procedure, $\textrm{Approx}(\cdot)$, returns within $O(d^{2}\log T)$ time under proper conditions. OQNS controls the number of exact Hessian-inverse evaluations to $O(\sqrt{T\log T})$, leading to a total runtime of $O(d^{2}T\log T+d^{\omega}\sqrt{T\log T})$, while achieving the asymptotical optimal regret $O(d\log T)$.
@@
However, their improvement comes at the cost of introducing large constant factors and abandoning the OMD framework of ONS. Specifically, when $\alpha\geq\frac{1}{DG}$, the leading $\log T$ term of ONS carries coefficient $DGd$ in Eq.˜1, whereas OQNS’s leading coefficient is $\frac{51}{2}DGd$ in Eq.˜5. Moreover, the log-barrier limits the applicability of OQNS, which will be dicussed in Section˜5.
Algorithm 1 ONS (Hazan et al., 2007)
@@
1: Initialize $\gamma_{0}=\frac{1}{2}\min\left\{\frac{1}{DG},\alpha\right\}$, $A_{0}=\epsilon I$, $\mathbf{x}_{1}=\bm{0}$.
@@
0: preconditioner coefficient $\epsilon$, hysteresis coefficient $k$.
1: Initialize $\gamma=\frac{1}{2}\min\left\{\frac{2}{(k+1)DG},\alpha\right\}$, $A_{0}=\epsilon I$, $\mathbf{x}_{1}=\bm{0}$.
@@
6:  $\mathbf{x}_{t+1}=\begin{cases}\widehat{\mathbf{x}}_{t+1}&\text{if}~\left\|\widehat{\mathbf{x}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{A_{t}}_{\mathcal{X}}[\widehat{\mathbf{x}}_{t+1}]&\text{otherwise}\end{cases}$.
@@ 3 Our Algorithm: LightONS
In this spirit, we present our algorithm LightONS. In Section˜3.1, to illustrate the key idea, we introduce the core algorithm that only differs from ONS by one line of code. However, the core algorithm is essentially improper learning. To address this issue, we introduce the improper-to-proper conversion in Section˜3.2, which yields the complete version of LightONS. Next, in Section˜3.3, we provide efficient implementations of LightONS. Finally, in Section˜3.4, we validate the empirical superiority of our method, which corroborates its theoretical guarantees.
3.1 Amortizing Projections with Hysteresis
The key idea of LightONS is to amortize the costly Mahalanobis projections with a hysteresis mechanism and enforce feasibility with computationally cheap Euclidean projections.
To illustrate this idea, we first introduce the core algorithm LightONS.Core in Algorithm˜2. As shown in Line 6, the proposed LightONS.Core performs a projection only when the decision moves outside an expanded domain $\widetilde{\mathcal{X}}\subseteq\mathbb{R}^{d}$ defined as
$\widetilde{\mathcal{X}}\triangleq\mathcal{B}(kD/2)=\{\mathbf{x}\in\mathbb{R}^{d}\,|\,\left\|\mathbf{x}\right\|_{2}\leq kD/2\},$ (8)
where $k>1$ is the hysteresis coefficient. Essentially, $\widetilde{\mathcal{X}}$ is the Euclidean ball of radius $kD/2$ centered at the origin, obtained by scaling by a factor of $k$ the minimal ball that contains $\mathcal{X}$. The key difference between LightONS.Core and the standard ONS algorithm (Hazan et al., 2007) is that ONS projects onto $\mathcal{X}$ immediately once the decision leaves $\mathcal{X}$, whereas LightONS.Core delays projection and continues updating outside $\mathcal{X}$, projecting back onto $\mathcal{X}$ only when the decision exits the expanded domain $\widetilde{\mathcal{X}}$.
Clearly, allowing a sufficiently large $k$ significantly improves projection efficiency, since projections are invoked less frequently. The following lemma precisely characterizes the relationship between the hysteresis coefficient $k$ and the number of Mahalanobis projections performed by the core algorithm. The proof of this lemma is provided in Section˜B.1.
@@ 3.1 Amortizing Projections with Hysteresis
Under Assumption˜1, and that for any $t\in[T]$ the loss function $f_{t}:\widetilde{\mathcal{X}}\to\mathbb{R}$ is $\alpha$-exp-concave, differentiable and $G$-Lipschitz on $\widetilde{\mathcal{X}}$, let $N$ denote the number of Mahalanobis projections in LightONS.Core (Algorithm˜2) over $T$ rounds, then
@@
On the other hand, because the proposed hysteresis mechanism alters the projection rule, it is essential to verify that the regret guarantee is preserved. In particular, increasing $k$ degrades the curvature parameter $\gamma$, leading to weaker regret bounds. In the extreme case $k$ approaches infinity, $\gamma$ collapses to zero, and exp-concavity degenerates to mere convexity. The following theorem establishes the relationship between the hysteresis coefficient $k$ and the regret bound of LightONS.Core. The proof of this theorem is provided in Section˜B.2.
@@
Under Assumption˜1, and assuming that for any $t\in[T]$ the loss function $f_{t}:\widetilde{\mathcal{X}}\to\mathbb{R}$ is $\alpha$-exp-concave, differentiable and $G$-Lipschitz on $\widetilde{\mathcal{X}}$, LightONS.Core (Algorithm˜2) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{d}{2\gamma}\log\left(1+\frac{G^{2}}{d\epsilon}T\right)+\frac{\gamma\epsilon D^{2}}{8},$ (10)
where $\gamma=\frac{1}{2}\min\left\{\frac{2}{(k+1)DG},\alpha\right\}$. The total runtime of LightONS.Core over $T$ rounds is
$O\left(d^{2}T+(k-1)^{-1}\sqrt{dT/\epsilon}\cdot\texttt{MP}_{\mathcal{X}}\right).$ (11)
As $k$ increases, Eq.˜10 shows that the regret bound scales as $O(\frac{1}{\gamma})=O(k+1)$, whereas Eq.˜11 indicates that the total runtime spent on Mahalanobis projections decreases as $O(N)=O(\frac{1}{k-1})$. Choosing $k=2$ therefore yields a favorable regret-efficiency trade-off: the regret bound grows by at most a factor of $\frac{3}{2}$, while the number of projections is substantially reduced. This suggests that LightONS.Core already achieves a significant improvement over standard ONS in terms of computational efficiency without sacrificing regret guarantees. However, as discussed below, LightONS.Core actually suffers from an improper learning issue.
@@
However, our LightONS.Core algorithm falls in the scope of improper learning (Shalev-Shwartz and Ben-David, 2014), as the algorithm’s decisions $\mathbf{x}_{t}\in\widetilde{\mathcal{X}}\supset\mathcal{X}$ may reside beyond the domain while the comparator $\mathbf{u}\in\mathcal{X}$ is strictly constrained to the domain. The core algorithm requires additional assumptions that the Lipschitzness and exp-concavity of the loss functions extend to the enlarged domain $\widetilde{\mathcal{X}}$. Vitally, such additional assumptions suppress the theoretical performance limits of its proper counterparts. A notable illustration is online logistic regression, where an improper learner achieves regret bounds $O(d\log(GT))$ (Foster et al., 2018), while proper learners are limited to a regret lower bound $\Omega(d\mathrm{e}^{G}\log T)$ (Hazan et al., 2014).
In the next subsection, we build upon LightONS.Core to develop a learning algorithm for OXO that achieves efficient updates and optimal regret, while guaranteeing that all decisions remain within the domain $\mathcal{X}$.
@@ 3.2 The Improper-to-Proper Conversion
Based on techniques from parameter-free online learning (Cutkosky and Orabona, 2018; Cutkosky, 2020), we incorporate a domain conversion into LightONS.Core, which yields a fully proper algorithm LightONS. This incorporation is conceptually simple yet technically delicate, as it recovers the optimal regret of ONS while preserving the efficiency gains of LightONS.Core.
We present the complete version of LightONS in Algorithm˜3. The domain conversion works by constructing surrogate loss functions. The core algorithm inside LightONS interacts with the surrogate loss $g_{t}$ (instead of the true loss $f_{t}$) and generates the surrogate decision $\mathbf{y}_{t}\in\widetilde{\mathcal{X}}$ (instead of the true decision $\mathbf{x}_{t}\in\mathcal{X}$). The true decision $\mathbf{x}_{t}$ is then obtained as the Euclidean projection of $\mathbf{y}_{t}$ onto the domain $\mathcal{X}$. For LightONS, any surrogate loss satisfying the following condition accomplishes this conversion.
@@
For some $c_{f}\geq 1$ and $c_{g}\geq 1$, the surrogate loss function $g_{t}:\mathbb{R}^{d}\to\mathbb{R}$ satisfies that, for any $\mathbf{u}\in\mathcal{X}$, $\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}\leq c_{g}G$ and $\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\leq c_{f}\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})$.
We note that two types of such conversions have been proposed in the literature. The first, by Cutkosky and Orabona (2018), satisfies the condition with $c_{f}=2$ and $c_{g}=1$. Later, an improved conversion was proposed with $c_{f}=1$ and $c_{g}=1$ (Cutkosky, 2020). In this work, we use the latter, as it incurs smaller constants and leads to tighter regret bounds.
Lemma 3 (Theorem 2 of Cutkosky (2020)).
@@
then for any $\mathbf{u}\in\mathcal{X}$, $\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}\leq\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}\leq G$, and $\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\leq\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})$.
Algorithm 3 LightONS
0: preconditioner coefficient $\epsilon$, hysteresis coefficient $k$.
1: Initialize $\gamma^{\prime}=\frac{1}{2}\min\left\{\frac{1}{c_{f}c_{g}DG},\frac{4}{c_{f}c_{g}(k+1)DG},\alpha\right\}$; $A_{0}=\epsilon I$; $\mathbf{x}_{1}=\mathbf{y}_{1}=\bm{0}$.
2:for $t=1,\dots,T$ do
3:  Observe $\nabla f_{t}(\mathbf{x}_{t})$.
4:  Construct $\nabla g_{t}(\mathbf{y}_{t})$, where $g_{t}$ satisfies Condition˜1.
5:  $A_{t}=A_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}$.
6:  $\widehat{\mathbf{y}}_{t+1}=\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})$.
7:  $\mathbf{y}_{t+1}=\begin{cases}\widehat{\mathbf{y}}_{t+1}&\text{if}~\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{A_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{y}}_{t+1}]&\text{otherwise}\end{cases}$.
8:  $\mathbf{x}_{t+1}=\Pi_{\mathcal{X}}[\mathbf{y}_{t+1}]$.
9:end for
We remark that the computational overhead of this conversion is negligible. Constructing the surrogate gradient $\nabla g_{t}(\mathbf{y}_{t})$ as Eq.˜12b takes only $O(d)$ arithmetic operations per round apart from the Euclidean projection, veiled by the $O(d^{2})$ cost of updating and inverting the Hessin-related matrix $A_{t}$. It is worth noting that several prior works have also employed this domain-conversion technique to address projection-related issues, but for different purposes, such as in non-stationary online learning (Zhao et al., 2025) and universal online learning (Yang et al., 2024).
One caveat is that the surrogate loss function in Lemma˜3 is not exp-concave. Fortunately, it can still exploit the exp-concavity of the original loss function $f_{t}$ using the following lemma, which shows that the surrogate loss $g_{t}$ exhibits a curvature property similar to Lemma˜1. The proof of this lemma is provided in Section˜B.3.
@@
Under Assumptions˜1, 2, 3 and 1, for any $\mathbf{u}\in\mathcal{X}$,
@@
$\leq\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})-\frac{\gamma^{\prime}}{2}\left(\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})\right)^{2}$
where $\gamma^{\prime}=\frac{1}{2}\min\left\{\frac{1}{c_{f}c_{g}DG},\frac{4}{(k+1)c_{f}c_{g}DG},\alpha\right\}$ as in LightONS (Algorithm˜3).
Lemma˜4 shows that, when the hysteresis coefficient $k$ is small (specifically $k\leq 3$), the curvature parameter $\gamma^{\prime}$ is unimpaired relative to $\gamma_{0}$ in Lemma˜1, with the domain conversion in Lemma˜3 ensuring $c_{f}=c_{g}=1$.
The following theorem establishes the optimality and efficiency of our proposed LightONS. The proof of this theorem is provided in Section˜B.4.
@@
Under Assumptions˜1, 2 and 3, LightONS (Algorithm˜3) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{d}{2\gamma^{\prime}}\log\left(1+\frac{c_{g}^{2}G^{2}}{d\epsilon}T\right)+\frac{\gamma^{\prime}\epsilon D^{2}}{8},$ (14)
where $\gamma^{\prime}=\frac{1}{2}\min\left\{\frac{1}{c_{f}c_{g}DG},\frac{4}{c_{f}c_{g}(k+1)DG},\alpha\right\}$. The total runtime of LightONS over $T$ rounds is
$O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{2}\right)T+(k-1)^{-1}d^{\omega+0.5}\sqrt{T/\epsilon}\log T\right).$ (15)
The preceding theorem immediately yields the following corollary.
@@
Under Assumptions˜1, 2 and 3, LightONS (Algorithm˜3) with $k=2$ and the surrogate loss function from Lemma˜3 satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{d}{2\gamma_{0}}\log\left(1+\frac{G^{2}}{d\epsilon}T\right)+\frac{\gamma_{0}\epsilon D^{2}}{8},$ (16)
where $\gamma_{0}=\frac{1}{2}\min\left\{\frac{1}{DG},\alpha\right\}$. Furthermore, with $\epsilon=d\log T$, the total runtime of LightONS over $T$ rounds is:
$O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{2}\right)T+d^{\omega}\sqrt{T\log T}\right).$ (17)
@@
We note that the domain conversion slightly worsens the constants, as seen in the second inequality of Eq.˜13. Nonetheless, this degradation is dominated by the problem-dependent parameters of OXO, namely $T$, $d$, $D$, $G$, $\alpha$. In their dependence on these parameters, LightONS’s regret in Eq.˜16 exactly matches ONS’s regret in Eq.˜1.

### v2

@@ 2.2 Important Progress
$\Phi_{t}(\mathbf{x})$ $\triangleq\eta d\log\frac{1}{1-\left\|\mathbf{x}\right\|_{2}^{2}}+\frac{d+\eta}{2}\left\|\mathbf{x}\right\|_{2}^{2}+\sum_{s=1}^{t}\left(\nabla f_{s}(\mathbf{x}_{s})^{\mathsf{T}}\mathbf{x}+\frac{\gamma}{2}\left(\nabla f_{s}(\mathbf{x}_{s})^{\mathsf{T}}(\mathbf{x}-\mathbf{x}_{s})\right)^{2}\right).$
@@
Evaluating the Hessian-inverse still takes $O(d^{\omega})$ time and OQNS mitigates this issue by approximating the Hessian-inverse gradient product with incremental updates. The approximation procedure, $\textrm{Approx}(\cdot)$, returns within $O(d^{2}\log T)$ time under proper conditions. OQNS controls the number of exact Hessian-inverse evaluations to $O(\sqrt{T\log T})$, leading to a total runtime of $O(d^{2}T\log T+d^{\omega}\sqrt{T\log T})$, while achieving the optimal regret $O(d\log T)$.
@@
However, their improvement introduces large constant factors. Specifically, when $\alpha\geq\frac{1}{DG}$, the leading $\log T$ term of ONS carries coefficient $DGd$ in Eq.˜1a, whereas OQNS’s leading coefficient is $\frac{51}{2}DGd$ in Eq.˜3a. Moreover, OQNS departs from the OMD framework of ONS, limiting its adaptability to broader scenarios, which will be discussed in Section˜5.
Algorithm 1 ONS [Hazan et al., 2007]
@@
1: Initialize $\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$, $A_{0}=\epsilon I$, $\mathbf{x}_{1}=\bm{0}$.
@@
0: preconditioner coef. $\epsilon$, deferral coef. $k$.
1: Initialize $\gamma=\frac{1}{2}\min\{\frac{2}{(k+1)DG},\alpha\}$, $A_{0}=\epsilon I$, $\mathbf{x}_{1}=\bm{0}$.
@@
6:  $\mathbf{x}_{t+1}=\begin{cases}\widehat{\mathbf{x}}_{t+1}&\text{if}\penalty 10000\ \left\|\widehat{\mathbf{x}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{A_{t}}_{\mathcal{X}}[\widehat{\mathbf{x}}_{t+1}]&\text{otherwise}\end{cases}$.
@@ 3 Our Algorithm: LightONS
In this spirit, we present our algorithm LightONS. In Section˜3.1, to illustrate the key idea, we introduce the core algorithm that only differs from ONS by one line of code. However, the core algorithm is essentially improper learning. To address this issue, we introduce the improper-to-proper conversion in Section˜3.2, which yields the complete version of LightONS.
Furthermore, we empirically validate the superiority of our method, which corroborates theoretical guarantees. Details of the experiments are deferred to Appendix˜A.
3.1 Amortizing Deferred Projections
We first introduce the LightONS.Core in Algorithm˜2, which amortizes the costly Mahalanobis projections with a deferred-projection mechanism. LightONS.Core only differs from ONS by one line of code, as shown in Line 6 of Algorithms˜1 and 2. While ONS projects onto $\mathcal{X}$ immediately when the decision exits $\mathcal{X}$, LightONS.Core continues to update without projection outside $\mathcal{X}$, and projects only when the decision exits an expanded domain $\widetilde{\mathcal{X}}_{k}\subseteq\mathbb{R}^{d}$ defined as
$\widetilde{\mathcal{X}}_{k}\triangleq\mathcal{B}(kD/2)=\{\mathbf{x}\in\mathbb{R}^{d}\,|\,\left\|\mathbf{x}\right\|_{2}\leq kD/2\},$ (5)
where $k>1$ is the deferral coefficient. Essentially, $\widetilde{\mathcal{X}}_{k}$ is obtained by scaling by a factor of $k$ the minimal ball that contains $\mathcal{X}$.
Clearly, a larger $k$ implies lower runtime, as it invokes fewer Mahalanobis projections. The following lemma quantifies the relationship between the deferral coefficient $k$ and the number of Mahalanobis projections $N$. Its proof is deferred to Section˜C.1.
@@ 3.1 Amortizing Deferred Projections
Under Assumption˜1, and that for any $t\in[T]$ the loss function $f_{t}$ is $\alpha$-exp-concave, differentiable and $G$-Lipschitz on $\widetilde{\mathcal{X}}_{k}$, let $N$ denote the number of Mahalanobis projections in LightONS.Core (Algorithm˜2) over $T$ rounds, then
@@
On the other hand, a larger $k$ exacerbates the deviation of LightONS.Core from ONS, potentially harming the regret guarantee. In particular, increasing $k$ degrades the curvature parameter $\gamma$, jeopardizing the curvature benefit of exp-concavity. In the extreme case where $k$ approaches infinity, $\gamma$ collapses to zero, and exp-concavity degenerates to mere convexity. The following theorem establishes the theoretical guarantees of LightONS.Core, revealing the trade-off between regret and runtime induced by the deferral coefficient $k$. Its proof is deferred to Section˜C.2.
@@
Under Assumption˜1, and that for any $t\in[T]$ the loss function $f_{t}$ is $\alpha$-exp-concave, differentiable and $G$-Lipschitz on $\widetilde{\mathcal{X}}_{k}$, with $\gamma=\frac{1}{2}\min\{\frac{2}{(k+1)DG},\alpha\}$, LightONS.Core (Algorithm˜2) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\frac{d}{2\gamma}\log\left(1+\frac{G^{2}}{d\epsilon}T\right)+\frac{\gamma\epsilon D^{2}}{8},$ (7a)
Runtime $\leq O\left(d^{2}T+(k-1)^{-1}\sqrt{dT/\epsilon}\cdot\texttt{MP}_{\mathcal{X}}\right).$ (7b)
As $k$ increases, Eq.˜7a shows that the regret bound scales as $O(\frac{1}{\gamma})=O(k+1)$, whereas Eq.˜7b indicates that the total runtime spent on Mahalanobis projections decreases as $O(N)=O(\frac{1}{k-1})$. With $k=2$, LightONS.Core already achieves a significant runtime improvement over ONS without sacrificing the optimal regret. Specifically, the regret bound grows by at most a factor of $\frac{3}{2}$, while the number of projections is substantially reduced from $O(T)$ to $O(\sqrt{T})$.
@@
However, LightONS.Core falls in the scope of improper learning [Shalev-Shwartz and Ben-David, 2014], as the algorithm’s decisions $\mathbf{x}_{t}\in\widetilde{\mathcal{X}}_{k}\supset\mathcal{X}$ may reside beyond the domain while the comparator $\mathbf{u}\in\mathcal{X}$ is strictly constrained to the domain. Moreover, LightONS.Core requires additional assumptions that the Lipschitzness and exp-concavity of the loss functions extend to the expanded domain $\widetilde{\mathcal{X}}_{k}$. Vitally, improper learning suppresses the theoretical performance limits of its proper counterparts. A notable illustration is online logistic regression, where an improper learner achieves the regret bound of $O(d\log(GT))$ [Foster et al., 2018], while proper learners are limited to a regret lower bound of $\Omega(d\mathrm{e}^{G}\log T)$ [Hazan et al., 2014].
@@ 3.2 The Improper-to-Proper Conversion
Based on techniques from parameter-free online learning [Cutkosky and Orabona, 2018; Cutkosky, 2020], we incorporate a domain conversion into LightONS.Core, yielding LightONS, a fully proper algorithm. This incorporation is conceptually straightforward yet technically subtle, as it inherits the favorable regret-runtime trade-off of LightONS.Core while ensuring proper learning over the domain $\mathcal{X}$. LightONS is summarized in Algorithm˜3. A detailed discussion of the practical implementation of LightONS is deferred to Section˜B.3.
Algorithm 3 LightONS
0: preconditioner coefficient $\epsilon$, deferral coefficient $k$.
1: Initialize $\gamma^{\prime}=\frac{1}{2}\min\{\frac{1}{c_{f}c_{g}DG},\frac{4}{c_{f}c_{g}(k+1)DG},\alpha\}$, $A_{0}=\epsilon I$, $\mathbf{x}_{1}=\mathbf{y}_{1}=\bm{0}$.
2:for $t=1,\dots,T$ do
3:  Observe $\nabla f_{t}(\mathbf{x}_{t})$; and construct $\nabla g_{t}(\mathbf{y}_{t})$ satisfying Condition˜1.
4:  $A_{t}=A_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}$.
5:  $\widehat{\mathbf{y}}_{t+1}=\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})$.
6:  $\mathbf{y}_{t+1}=\begin{cases}\widehat{\mathbf{y}}_{t+1}&\text{if}\penalty 10000\ \left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{A_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{y}}_{t+1}]&\text{otherwise}\end{cases}$.
7:  $\mathbf{x}_{t+1}=\Pi_{\mathcal{X}}[\mathbf{y}_{t+1}]$.
8:end for
The domain conversion proceeds by constructing surrogate loss functions and projecting surrogate decisions. At each round, LightONS constructs a surrogate loss $g_{t}:\mathbb{R}^{d}\to\mathbb{R}$ from the true loss $f_{t}:\mathcal{X}\to\mathbb{R}$ and supplies $g_{t}$ to to an underlying procedure of LightONS.Core; The underlying LightONS.Core outputs a surrogate decision $\mathbf{y}_{t}\in\widetilde{\mathcal{X}}_{k}$, which LightONS then maps to a (proper) true decision $\mathbf{x}_{t}\in\mathcal{X}$ via a Euclidean projection. For LightONS, any surrogate loss satisfying the condition below suffices to guarantee a valid conversion.
@@
For some $c_{f}\geq 1$ and $c_{g}\geq 1$, the surrogate loss function $g_{t}:\mathbb{R}^{d}\to\mathbb{R}$ satisfies that, for any $\mathbf{u}\in\mathcal{X}$, $\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}\leq c_{g}\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}$ and $\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\leq c_{f}\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})$.
We note that two such conversions have been proposed in the literature. The first, by Cutkosky and Orabona [2018], achieves $c_{f}=2$ and $c_{g}=1$. Later, an improved conversion by Cutkosky [2020] achieves $c_{f}=c_{g}=1$. We prefer the latter which yields smaller constants and tighter regret.
Lemma 3 (Theorem 2 of Cutkosky [2020]).
@@
then for any $\mathbf{u}\in\mathcal{X}$, $\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}\leq\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}$, and $\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\leq\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})$.
The computational overhead of this conversion is negligible. Constructing the surrogate gradient $\nabla g_{t}(\mathbf{y}_{t})$ as Eq.˜8b takes only $O(d)$ arithmetic operations per round apart from the Euclidean projection, veiled by the $O(d^{2})$ cost of updating and inverting the Hessian-related matrix $A_{t}$. It is worth noting that prior work has also employed this domain-conversion technique to address projection-related issues, but for different purposes, such as in non-stationary online learning [Zhao et al., 2025] and universal online learning [Yang et al., 2024].
One caveat is that the surrogate loss $g_{t}$ in Lemma˜3 is actually not exp-concave, seemingly precluding the curvature benefit. Fortunately, the following lemma shows that the surrogate loss $g_{t}$ inherits the curvature in a form closely mirroring Lemma˜1, with proof in Section˜C.3.
@@
Under Assumptions˜1, 2 and 3, with $\gamma^{\prime}=\frac{1}{2}\min\{\frac{1}{c_{f}c_{g}DG},\frac{4}{(k+1)c_{f}c_{g}DG},\alpha\}$, for any $\mathbf{u}\in\mathcal{X}$ and any surrogate loss function $g_{t}$ satisfying Condition˜1,
@@
$\leq\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})-\frac{\gamma^{\prime}}{2}\left(\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})\right)^{2}.$
We remark that, when the deferral coefficient $k\leq 3$ and $c_{f}=c_{g}=1$ (e.g., in Lemma˜3), the surrogate curvature parameter $\gamma^{\prime}$ in Lemma˜4 is unimpaired relative to the original $\gamma_{0}$ in Lemma˜1.
The following theorem establishes the theoretical guarantees of LightONS, and an immediate corollary specifies its regret and runtime by instantiating deferral coefficient $k=2$ and the surrogate loss in Lemma˜3. Its proof is deferred to Section˜C.4.
@@
Under Assumptions˜1, 2 and 3, with $\gamma^{\prime}=\frac{1}{2}\min\{\frac{1}{c_{f}c_{g}DG},\frac{4}{c_{f}c_{g}(k+1)DG},\alpha\}$, LightONS (Algorithm˜3) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\frac{d}{2\gamma^{\prime}}\log\left(1+\frac{c_{g}^{2}G^{2}}{d\epsilon}T\right)+\frac{\gamma^{\prime}\epsilon D^{2}}{8},$ (10a)
Runtime $\leq O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{2}\right)T+(k-1)^{-1}d^{\omega+0.5}\sqrt{T/\epsilon}\log T\right).$ (10b)
@@
Under Assumptions˜1, 2 and 3, by setting the surrogate loss function as Lemma˜3, $k=2$ and $\epsilon=d\log T$, with $\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$, LightONS (Algorithm˜3) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\frac{d}{2\gamma_{0}}\log\left(1+\frac{G^{2}}{d\epsilon}T\right)+\frac{\gamma_{0}\epsilon D^{2}}{8},$ (11a)
Runtime $\leq O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{2}\right)T+d^{\omega}\sqrt{T\log T}\right).$ (11b)
@@
LightONSachieves computational efficiency at the cost of a slightly worse regret than ONS, as seen in the gap of the second inequality in Eq.˜9. Nonetheless, this degradation is dominated by the problem-dependent parameters of OXO, namely $T$, $d$, $D$, $G$, $\alpha$. In their dependence on these parameters, LightONS’s regret in Eq.˜11a exactly matches ONS’s regret in Eq.˜1a.